OpenStudy (anonymous):

Solve the following system of equations: 5x+4y=25 3x+4y=15

OpenStudy (anonymous):

x=5;

OpenStudy (anonymous):

subtract the second equation from the first

OpenStudy (anonymous):

5x+4y - (3y+4y)=25 - 15 simplify and solve :-)

OpenStudy (anonymous):

no solution for y

OpenStudy (anonymous):

http://www.purplemath.com/modules/systlin1.htm

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

5x+4y=25 3x+4y=15 5x+4y -(3x+4y)=25-15 2x=10 so x=5 if x=5 then 5x+4y=25 yields, 5(5)+4y=25 so 25+4y=25 which yields y=0 so your solution is (5,0) meaning at the point (5,0) your two lines intersect